# Deployability-Centric Infrastructure-as-Code Generation: Fail, Learn, Refine, and Succeed through LLM-Empowered DevOps Simulation

### v2

@@ 1. Introduction
The digital transformation of enterprises has fundamentally altered how organizations manage and deploy computing infrastructure. As cloud adoption accelerates globally, with over 82% of enterprises now utilizing cloud services (rackspace-report), the complexity and scale of infrastructure management have increased exponentially. Infrastructure-as-Code (IaC) has emerged as the default standard for managing this complexity, enabling organizations to define, provision, and maintain cloud resources through programmable templates that ensure consistency, repeatability, and version control across deployments (kumara2021s; chinamanagonda2019automating; sokolowski2022infrastructure). This paradigm shift from manual infrastructure management to code-driven automation has become essential for modern DevOps practices (basher2019devops), supporting the rapid deployment cycles and scalability requirements of contemporary software systems.
@@
However, the widespread adoption of IaC faces significant barriers rooted in its inherent technical complexity and the specialized expertise it demands (cusumano2010cloud; qiu2023simplifying; iac_low). Creating IaC templates requires a deep understanding of cloud service architectures, resource interdependencies, and platform-specific constraints that extend far beyond traditional programming knowledge (basher2019devops). The steep learning curve of IaC languages creates substantial obstacles for development teams lacking dedicated DevOps expertise (kon2024iac; hashicorp2023). This expertise gap not only slows the development speed but also increases the risk of misconfigurations that lead to security vulnerabilities and service outages (nasiri2024towards; drosos2024your).
@@
Recent advances in Large Language Models (LLMs) present a promising opportunity to democratize IaC development by generating deployable infrastructure templates from natural language descriptions (chen2021evaluating; liu2023your; liao2024code; liu2025ai). An effective LLM-based IaC generator could potentially transform infrastructure automation by enabling developers to describe their infrastructure requirements in natural language and receive working deployable IaC in return. This capability would significantly accelerate development workflows in DevOps and reduce the expertise barrier for cloud adoption.
@@
However, previous research indicates that LLMs demonstrate considerably lower performance in IaC generation compared to general programming tasks (kon2024iac; ragothaman2024optimizing; han2024chase), with initial success rates as low as 19%. We notice that existing research (kon2024iac; ragothaman2024optimizing; palavalli2024using) suffers from a critical limitation: an overemphasis on syntactic correctness while neglecting the fundamental requirement of deployability. In operational contexts, deployability represents the primary measure of IaC template quality, as non-deployable templates provide no practical value regardless of syntactic correctness.
@@
To mitigate research gaps, we present DPIaC-Eval, a novel benchmark comprising 153 real-world infrastructure tasks across 58 unique AWS services, with expert-authored natural language requirements and verified deployable IaC template as reference templates. The formative study shows that mainstream LLMs can only achieve a merely 50% deployment success rate even in the simplest scenarios (with only one resource and one parameter) in its first attempt. This poor performance stems from fundamental challenges in IaC generation: unlike traditional programming where logic errors are often apparent, IaC’s declarative nature obscures issues until execution, requiring iterative testing and validation. Therefore, we propose IaCGen, a novel deployability-centric IaC generation framework. IaCGen is integrated through an iterative feedback mechanism across three deployability-oriented validation stages, format verification, syntax checking, and live deployment, which mirrors the real DevOps workflows. Such design enables models to learn from syntactic and deployment failures, progressively refining templates.
@@ 2.2. Difference Between IaC Tools
Terraform (TF) and CloudFormation (CF) represent two distinct types of IaC tools, each with unique format language, design philosophy, and operational characteristics. TF uses the HashiCorp Configuration Language (HCL), a declarative language designed with a provider-based architecture that enables multi-cloud deployments across AWS, Azure, Google Cloud, and numerous other platforms (pum2024cloudformation), while CF uses JSON or YAML to define resources in a hierarchical structure with deep integration into the AWS’s service ecosystem (hash-tf-cf-compare). Additionally, CF focuses exclusively on AWS, providing native integration with AWS services, security models, and deployment patterns, including features like AWS-specific intrinsic functions and cross-stack references (geeks-tf-cf-compare).
In addition to the syntactic validity, the program deployability fatally determines real-world IaC template utility. TF’s stateful architecture requires explicit state management and manual resource cleanup procedures (geeks-tf-cf-compare), making systematic deployability evaluation across hundreds of templates impractical. Also, TF’s lack of native rollback mechanisms means that failed deployments can leave partially created resources in indeterminate states, requiring manual cloud resources cleaning (tf-limitation-first; tf-limitation-second; hash-tf-cf-compare). These operational constraints might explain why existing TF-based benchmark limit evaluation to static analysis (kon2024iac), leaving the crucial deployability dimension unaddressed.
CF, as AWS’s native IaC language, provides unique methodological advantages that make comprehensive deployability testing feasible. When deploying a CF template, AWS automatically creates a stack to hold the collection of resources defined in the template (cf-stack). Users can create, update, and delete entire resource collections by manipulating stacks, enabling systematic removal of all created resources and restoration of the cloud environment to its initial state through simple stack deletion (cf-stack). Therefore, to explore the automated IaC template generation especially for deployability, we focus on the CloudFormation in this paper.
@@ 3. Benchmark: DPIaC-Eval
Most existing research in code generation has predominantly concentrated on general-purpose programming languages (liao2024code; gu2023llm; koziolek2024llm; coignion2024performance), often overlooking the distinct challenges inherent to IaC. Unlike conventional software programs, IaC templates must not only be syntactically correct but also semantically valid and deployable within specific cloud environments. Thus, to systematically evaluate the effectiveness of IaC template generation, we develop the first deployability-focused benchmark, DPIaC-Eval, comprising 153 diverse CloudFormation IaC templates.
@@ 3.1. Benchmark Construction
Template Sources. Templates are sourced from two primary channels to ensure diversity and quality: (1) official AWS documentation and sample libraries, including the AWS CloudFormation Sample Templates repository (aws_cfn_templates); (2) GitHub repositories that contain CloudFormation template which we find within the AWS Samples GitHub organization (aws_samples_github) and a dataset containing URLs of repositories that use CloudFormation templates (cirlan2024mining). Ethical considerations are enforced by confirming appropriate licensing (e.g., MIT, Apache 2.0) before inclusion, ensuring that all templates are legally permissible for research purposes.
Preprocessing.  Following an initial collection of approximately 900 templates, candidates undergo automated preprocessing through a multistage filtering funnel. First, size filtering removes templates that exceed 8,000 tokens or 600 lines, ensuring compatibility with the common LLM generation capacities, resulting in 850 templates. Next, syntax validation verifies templates for structural correctness using cfn-linter (cfn-linter), yielding 465 valid candidates. Templates then undergo deployment testing using the AWS SDK (boto3) with a sandbox AWS account configured with least-privilege IAM roles, reducing the dataset to 200 templates. Finally, a rectification process addresses common deployment failures, including cross-stack dependencies, parameter value completion, missing VPC resources, outdated value references, and dependency ordering adjustments. After rectification, our final dataset comprises 153 high-quality, deployable CloudFormation templates that serve as the foundation for our subsequent IaC generation experiments.
@@ 3.2. Benchmark Characteristics
Each entry in the benchmark contains two essential elements: (1) a natural language prompt describing the infrastructure requirements and (2) the reference CloudFormation template file path. The benchmark comprises 58 commonly used AWS cloud infrastructure services (drosos2024your; jain2023skyplane; stoica2021cloud). Fig. 3(a) illustrates the 20 most frequent services in our dataset, with frequency represented by the number of templates in which each service appears. The distribution spans the full stack of infrastructure components typically required in modern cloud environments: (1) Compute and serverless: EC2, Lambda, and AutoScaling; (2) Data storage and management: S3, DynamoDB, RDS, and Elastic File System (EFS); (3) Networking infrastructure: Virtual Private Cloud (VPC) and API Gateway; (4) Security and management: Identity and Access Management (IAM), CloudWatch, Logs, and Systems Manager (SSM), Key Management Service (KMS), and SecretsManager; (5) Event processing and messaging: EventBridge (Events), Simple Queue Service (SQS), Pipes, and Simple Notification Service (SNS); (6) Data stream processing: Kinesis. This comprehensive coverage ensures our benchmark effectively evaluates LLMs’ capabilities across the diverse range of services essential for real-world cloud infrastructure deployment. The service distribution shown in
@@ 5 $\geq 200$ $\geq 12$ $\geq 14$
To facilitate systematic evaluation, we classify templates into five difficulty levels, building upon the design established by IaC-Eval (kon2024iac). Acknowledging the inherently subjective nature of drawing boundaries for different difficulty levels, the classification is based on objective metrics including the line of code (LoC) in the IaC template, the number of resources required, and the number of configured parameters. The criteria of difficulty level (1-5) are shown in Table 1. As shown in Fig. 3(b), the benchmark demonstrates a balanced distribution across difficulty levels. This distribution and range ensure the benchmark effectively challenges automated generation methods in various IaC template scenarios while maintaining real-world relevance.
@@ 3.3. Comparison with IaC-Eval Benchmark
Both DPIaC-Eval and IaC-Eval (kon2024iac) cover around 60 AWS services, though each includes unique services not found in the other. DPIaC-Eval presents greater challenges than existing IaC benchmarks, with longer average line of code (155 vs. 42) and more resources per template (7 vs. 4). More importantly, DPIaC-Eval provides comprehensive coverage in four critical dimensions: syntax correctness, deployability, user intent matching, and security. In contrast, IaC-Eval supports only syntax validation and user intent matching. This multi-dimensional validation approach enables our benchmark to identify critical gaps in generated IaC template that remain undetected in syntax-only evaluations. As we report in Section 6, the deployment validation reveals 42.7% of the syntactically correct templates fail during deployment, highlighting the importance of testing beyond mere syntax. Similarly, the security validation identifies only 9.0% of the deployable templates meet security compliance standards, exposing a critical shortcoming that is overlooked by IaC-Eval.
@@ 4. Framework
LLMs have achieved remarkable success in code generation tasks (lachaux2020unsupervised; roziere2021leveraging; si2024solution; pan2024large), demonstrating state-of-the-art performance across programming languages. However, their application to IaC generation remains limited and problematic. Current approaches (kon2024iac; ragothaman2024optimizing; pujar2023automated) treat IaC templates as conventional programs, focusing primarily on syntactic correctness while neglecting their continuous integration and deployment (CI/CD) nature (srivatsa2024survey; iac-ci-cd). Consequently, generated templates often fail during actual deployment despite being syntactically valid and resulting in significantly lower accuracy rates compared to general programming tasks (GPT 4 achieves 19% in IaC-Eval and 83% in HumanEval). Our framework, IaCGen, addresses this limitation through a deployment-driven approach that enables LLMs to learn from deployment failures and continuously refine templates until achieving successful deployment. As shown in Fig. 5, IaCGen consists of two modules, Iterative IaC Generation and IaC Deployment, integrated through a tiered feedback mechanism that mirrors real DevOps workflows. The complete prompts and other reproducible details mentioned in this section are available in our code repository.
@@ 4.2. IaC Deployment
This module takes the LLM-generated IaC template as input and validates its soundness through a three-stage deployment. Similarly to real-world testing, this process consists of sequential validations, from format verification to syntax checking to live deployment, with LLM iteratively refining the template based on error messages (iac-best-practices; vessey1985expertise; debugging-web). If an error occurs at any stage, feedback is generated and returned to the Iterative IaC Generation module. If the template passes all three validation stages, the framework will end with the current IaC template as the final output (step 8).
Format Verification. Most IaC solutions use declarative definition files to specify infrastructure configurations. The syntax for describing IaC usually depends on the requirements of the target platform, with common formats including YAML and JSON. The first validation stage verifies whether the LLM-generated IaC template conforms to the correct format specification. As AWS CloudFormation templates written in YAML, we employ yamllint (yaml-linter) to validate basic structural correctness, including proper indentation, key-value pair formatting, and quotation usage.
Syntax Checking. After confirming proper formatting, the second validation stage uses a static analysis tool, cfn-linter (cfn-linter), to verify if the code is correct in CloudFormation syntax. Although most existing LLM-based IaC frameworks stop after this stage, our approach extends validation to actual deployment testing, which addresses a critical gap in operational viability assessment.
Live Deployment. IaC shares a critical characteristic with traditional programming: code that passes compilation or syntax validation does not necessarily execute successfully at runtime. However, this challenge is particularly acute with IaC because of its unique properties. Most IaC languages are declarative rather than imperative, specifying the desired end state rather than procedural steps. This fundamental difference creates a validation gap that static analysis alone cannot bridge (guerriero2019adoption). Cloud environments impose numerous constraints that syntax validators cannot fully capture: regional service availability variations, deprecated resource identifiers, and service interdependencies. For example, a syntactically perfect Terraform template referencing an outdated Amazon Machine Image ID will pass linting but fail during actual provisioning. Thus, we employ live deployment validation with the AWS SDK boto3 client, the official toolkit developed by AWS.

### v3

@@ 1. Introduction
The digital transformation of enterprises has fundamentally altered how organizations manage and deploy computing infrastructure. As cloud adoption accelerates globally, with over 82% of enterprises now utilizing cloud services (Rackspace, 2025), the complexity and scale of infrastructure management have increased exponentially. Infrastructure-as-Code (IaC) has emerged as the default standard for managing this complexity, enabling organizations to define, provision, and maintain cloud resources through programmable templates that ensure consistency, repeatability, and version control across deployments (Kumara et al., 2021; Chinamanagonda, 2019; Sokolowski, 2022). This paradigm shift from manual infrastructure management to code-driven automation has become essential for modern DevOps practices (Basher, 2019), supporting the rapid deployment cycles and scalability requirements of contemporary software systems.
@@
However, the widespread adoption of IaC faces significant barriers rooted in its inherent technical complexity and the specialized expertise it demands (Cusumano, 2010; Qiu et al., 2023; Styra, 2023). Creating IaC templates requires a deep understanding of cloud service architectures, resource interdependencies, and platform-specific constraints that extend far beyond general programming knowledge (Basher, 2019; Li et al., 2024). The steep learning curve of IaC languages creates substantial obstacles for development teams lacking dedicated DevOps expertise (Kon et al., 2024; Hashicorp, 2023). This expertise gap not only slows the development speed but also increases the risk of misconfigurations that lead to security vulnerabilities and service outages (Nasiri et al., 2024; Drosos et al., 2024).
@@
Recent advances in Large Language Models (LLMs) present a promising opportunity to democratize IaC development by generating deployable infrastructure templates from natural language descriptions (Chen et al., 2021; Liu et al., 2023; Liao et al., 2024; Liu et al., 2025; Liao et al., 2025; Zhou et al., 2025). An effective LLM-based IaC generator could potentially transform infrastructure automation by enabling developers to describe their infrastructure requirements in natural language and receive working deployable IaC in return. This capability would significantly accelerate development workflows in DevOps and reduce the expertise barrier for cloud adoption.
@@
However, previous research indicates that LLMs demonstrate considerably lower performance in IaC generation compared to general programming tasks (Kon et al., 2024; Ragothaman and Udayakumar, 2024; Han et al., 2024), with initial success rates as low as 19%. We notice that existing research (Kon et al., 2024; Ragothaman and Udayakumar, 2024; Palavalli and Santolucito, 2024) suffers from a critical limitation: an overemphasis on syntactic correctness while neglecting the fundamental requirement of deployability. In operational contexts, deployability represents the primary measure of IaC template quality, as non-deployable templates provide no practical value regardless of syntactic correctness.
@@
To mitigate research gaps, we present DPIaC-Eval, a novel benchmark comprising 153 real-world infrastructure tasks across 58 unique AWS services, with expert-authored natural language requirements and verified deployable IaC template as reference templates. The formative study shows that mainstream LLMs can only achieve a merely 50% deployment success rate even in the simplest scenarios (with only one resource and one parameter) in its first attempt. This poor performance stems from fundamental challenges in IaC generation: unlike traditional programming where logic errors are often apparent (Sun et al., 2021), IaC’s declarative nature obscures issues until execution, requiring iterative testing and validation (Sun et al., 2023a, b, 2022). Therefore, we propose IaCGen, a novel deployability-centric IaC generation framework. IaCGen is integrated through an iterative feedback mechanism across three deployability-oriented validation stages, format verification, syntax checking, and live deployment, which mirrors the real DevOps workflows. Such design enables models to learn from syntactic and deployment failures, progressively refining templates.
@@ 2.2. Difference Between IaC Tools
Terraform (TF) and CloudFormation (CF) represent two distinct types of IaC tools, each with unique format language, design philosophy, and operational characteristics. TF uses the HashiCorp Configuration Language (HCL), a declarative language designed with a provider-based architecture that enables multi-cloud deployments across AWS, Azure, Google Cloud, and numerous other platforms (Pum and George, 2024), while CF uses JSON or YAML to define resources in a hierarchical structure with deep integration into the AWS’s service ecosystem (Hashicorp, 2024). Additionally, CF focuses exclusively on AWS, providing native integration with AWS services, security models, and deployment patterns, including features like AWS-specific intrinsic functions and cross-stack references (GeeksForGeeks, 2024).
In addition to the syntactic validity, the program deployability fatally determines real-world IaC template utility. TF’s stateful architecture requires explicit state management and manual resource cleanup procedures (GeeksForGeeks, 2024), making systematic deployability evaluation across hundreds of templates impractical. Also, TF’s lack of native rollback mechanisms means that failed deployments can leave partially created resources in indeterminate states, requiring manual cloud resources cleaning (Sean, 2022; HashiCorp, 2023; Hashicorp, 2024). These operational constraints might explain why existing TF-based benchmark limit evaluation to static analysis (Kon et al., 2024), leaving the crucial deployability dimension unaddressed.
CF, as AWS’s native IaC language, provides unique methodological advantages that make comprehensive deployability testing feasible. When deploying a CF template, AWS automatically creates a stack to hold the collection of resources defined in the template (AWS, 2025c). Users can create, update, and delete entire resource collections by manipulating stacks, enabling systematic removal of all created resources and restoration of the cloud environment to its initial state through simple stack deletion (AWS, 2025c). Therefore, to explore the automated IaC template generation especially for deployability, we focus on the CloudFormation in this paper.
@@ 3. Benchmark: DPIaC-Eval
Most existing research in code generation has predominantly concentrated on general-purpose programming languages (Liao et al., 2024; Gu, 2023; Koziolek et al., 2024; Coignion et al., 2024), often overlooking the distinct challenges inherent to IaC. Unlike conventional software programs, IaC templates must not only be syntactically correct but also semantically valid and deployable within specific cloud environments. Thus, to systematically evaluate the effectiveness of IaC template generation, we develop the first deployability-focused benchmark, DPIaC-Eval, comprising 153 diverse CloudFormation IaC templates.
@@ 3.1. Benchmark Construction
Template Sources. Templates are sourced from two primary channels to ensure diversity and quality: (1) official AWS documentation and sample libraries, including the AWS CloudFormation Sample Templates repository (1); (2) GitHub repositories that contain CloudFormation template which we find within the AWS Samples GitHub organization (Samples, 2025) and a dataset containing URLs of repositories that use CloudFormation templates (Cirlan, 2024). Ethical considerations are enforced by confirming appropriate licensing (e.g., MIT, Apache 2.0) before inclusion, ensuring that all templates are legally permissible for research purposes.
Preprocessing.  Following an initial collection of approximately 900 templates, candidates undergo automated preprocessing through a multistage filtering funnel. First, size filtering removes templates that exceed 8,000 tokens or 600 lines, ensuring compatibility with the common LLM generation capacities, resulting in 850 templates. Next, syntax validation verifies templates for structural correctness using cfn-linter (Kevin et al., 2025), yielding 465 valid candidates. Templates then undergo deployment testing using the AWS SDK (boto3) with a sandbox AWS account configured with least-privilege IAM roles, reducing the dataset to 200 templates. Finally, a rectification process addresses common deployment failures, including cross-stack dependencies, parameter value completion, missing VPC resources, outdated value references, and dependency ordering adjustments. After rectification, our final dataset comprises 153 high-quality, deployable CloudFormation templates that serve as the foundation for our subsequent IaC generation experiments.
@@ 3.2. Benchmark Characteristics
Each entry in the benchmark contains two essential elements: (1) a natural language prompt describing the infrastructure requirements and (2) the reference CloudFormation template file path. The benchmark comprises 58 commonly used AWS cloud infrastructure services (Drosos et al., 2024; Jain et al., 2023; Stoica and Shenker, 2021). Fig. 3(a) illustrates the 20 most frequent services in our dataset, with frequency represented by the number of templates in which each service appears. The distribution spans the full stack of infrastructure components typically required in modern cloud environments: (1) Compute and serverless: EC2, Lambda, and AutoScaling; (2) Data storage and management: S3, DynamoDB, RDS, and Elastic File System (EFS); (3) Networking infrastructure: Virtual Private Cloud (VPC) and API Gateway; (4) Security and management: Identity and Access Management (IAM), CloudWatch, Logs, and Systems Manager (SSM), Key Management Service (KMS), and SecretsManager; (5) Event processing and messaging: EventBridge (Events), Simple Queue Service (SQS), Pipes, and Simple Notification Service (SNS); (6) Data stream processing: Kinesis. This comprehensive coverage ensures our benchmark effectively evaluates LLMs’ capabilities across the diverse range of services essential for real-world cloud infrastructure deployment. The service distribution shown in
@@ 5 $\geq 200$ $\geq 12$ $\geq 14$
To facilitate systematic evaluation, we classify templates into five difficulty levels, building upon the design established by IaC-Eval (Kon et al., 2024). Acknowledging the inherently subjective nature of drawing boundaries for different difficulty levels, the classification is based on objective metrics including the line of code (LoC) in the IaC template, the number of resources required, and the number of configured parameters. The criteria of difficulty level (1-5) are shown in Table 1. As shown in Fig. 3(b), the benchmark demonstrates a balanced distribution across difficulty levels. This distribution and range ensure the benchmark effectively challenges automated generation methods in various IaC template scenarios while maintaining real-world relevance.
@@ 3.3. Comparison with IaC-Eval Benchmark
Both DPIaC-Eval and IaC-Eval (Kon et al., 2024) cover around 60 AWS services, though each includes unique services not found in the other. DPIaC-Eval presents greater challenges than existing IaC benchmarks, with longer average line of code (155 vs. 42) and more resources per template (7 vs. 4). More importantly, DPIaC-Eval provides comprehensive coverage in four critical dimensions: syntax correctness, deployability, user intent matching, and security. In contrast, IaC-Eval supports only syntax validation and user intent matching. This multi-dimensional validation approach enables our benchmark to identify critical gaps in generated IaC template that remain undetected in syntax-only evaluations. As we report in Section 6, the deployment validation reveals 42.7% of the syntactically correct templates fail during deployment, highlighting the importance of testing beyond mere syntax. Similarly, the security validation identifies only 9.0% of the deployable templates meet security compliance standards, exposing a critical shortcoming that is overlooked by IaC-Eval.
@@ 4. Framework
LLMs have achieved remarkable success in code generation tasks (Lachaux et al., 2020; Roziere et al., 2021; Si et al., 2024; Pan et al., 2024a), demonstrating state-of-the-art performance across programming languages. However, their application to IaC generation remains limited and problematic. Current approaches (Kon et al., 2024; Ragothaman and Udayakumar, 2024; Pujar et al., 2023) treat IaC templates as conventional programs, focusing primarily on syntactic correctness while neglecting their continuous integration and deployment (CI/CD) nature (Srivatsa et al., 2024; Azure, 2023; Liu et al., 2022). Consequently, generated templates often fail during actual deployment despite being syntactically valid and resulting in significantly lower accuracy rates compared to general programming tasks (GPT 4 achieves 19% in IaC-Eval and 83% in HumanEval). Our framework, IaCGen, addresses this limitation through a deployment-driven approach that enables LLMs to learn from deployment failures and continuously refine templates until achieving successful deployment. As shown in Fig. 5, IaCGen consists of two modules, Iterative IaC Generation and IaC Deployment, integrated through a tiered feedback mechanism that mirrors real DevOps workflows. The complete prompts and other reproducible details mentioned in this section are available in our code repository.
@@ 4.2. IaC Deployment
This module takes the LLM-generated IaC template as input and validates its soundness through a three-stage deployment. Similarly to real-world testing, this process consists of sequential validations, from format verification to syntax checking to live deployment, with LLM iteratively refining the template based on error messages (Iranpour, 2024; Vessey, 1985; saloman, 2023). If an error occurs at any stage, feedback is generated and returned to the Iterative IaC Generation module. If the template passes all three validation stages, the framework will end with the current IaC template as the final output (step 8).
Format Verification. Most IaC solutions use declarative definition files to specify infrastructure configurations. The syntax for describing IaC usually depends on the requirements of the target platform, with common formats including YAML and JSON. The first validation stage verifies whether the LLM-generated IaC template conforms to the correct format specification. As AWS CloudFormation templates written in YAML, we employ yamllint (et al., 2023) to validate basic structural correctness, including proper indentation, key-value pair formatting, and quotation usage.
Syntax Checking. After confirming proper formatting, the second validation stage uses a static analysis tool, cfn-linter (Kevin et al., 2025), to verify if the code is correct in CloudFormation syntax. Although most existing LLM-based IaC frameworks stop after this stage, our approach extends validation to actual deployment testing, which addresses a critical gap in operational viability assessment.
Live Deployment. IaC shares a critical characteristic with traditional programming: code that passes compilation or syntax validation does not necessarily execute successfully at runtime. However, this challenge is particularly acute with IaC because of its unique properties. Most IaC languages are declarative rather than imperative, specifying the desired end state rather than procedural steps. This fundamental difference creates a validation gap that static analysis alone cannot bridge (Guerriero et al., 2019). Cloud environments impose numerous constraints that syntax validators cannot fully capture: regional service availability variations, deprecated resource identifiers, and service interdependencies. For example, a syntactically perfect Terraform template referencing an outdated Amazon Machine Image ID will pass linting but fail during actual provisioning. Thus, we employ live deployment validation with the AWS SDK boto3 client, the official toolkit developed by AWS.